1) Instead of using zinc, a student accidentally used a 1 M solution of silver (Ag). Based on the information in the table below, what would be the theoretical cell voltage, Eocell, when it is used in a galvanic cell with copper?
Answer #1

now new reaction which will occur

2Ag+(aq) + 2e = 2Ag(s) E*= 0.80 V

Cu(s) = Cu2+(aq) + 2e E* = -0.34V

add both reactions

Cu(s) + 2Ag+(aq) = Cu2+(aq) + 2Ag(s)

E*(cell) = 0.80 + (-0.34)

= 0.46 V

any query please comment

if satisfied please rate it thnaks